OpenStudy (anonymous):

the quantities s and t are positive and are related by the equation s=k/t, where k is a constant. if the value of S increased by 50 percent, then the value of T decreases by what percent???

OpenStudy (anonymous):

is this a ratio ques??

OpenStudy (dumbcow):

well if you solve for t t = k/s Increasing by 50% is equivalent to multiplying by 1.5 t = k/(1.5s) 1/1.5 = 0.667 t = 0.667(k/s) t decreased by 33.3%

OpenStudy (anonymous):

well how did you get t=k/s

OpenStudy (dumbcow):

solve for t s = k/t multiply by t on both sides t*s = k divide by s on both sides t = k/s

OpenStudy (anonymous):

how did you conclude that increasing by 50% is equivalent to multip by 1.5

OpenStudy (dumbcow):

50% = 0.5 s + 50%s = s+0.5s = s(1+0.5) = 1.5s

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

so whenever something is increased i add the variable +.05 to x(1+.05)...also what would I do for a decrease ques??

OpenStudy (dumbcow):

if something decreases by say 30% x - 0.3x = x(1-.3) = 0.7x

OpenStudy (anonymous):

ok..cool thanks

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

why did you 1/1.5?? becuz k or 1k/1.5s??

OpenStudy (anonymous):

ok ok i got it

OpenStudy (anonymous):

but did you convert .667 into 33.1/3 %

OpenStudy (dumbcow):

got it? t = k/s the 1/1.5* (k/s) is the new value of t t is now 66.6% of its old value 100% - 66.6% = 33.3% so it decreased by 33.3%

OpenStudy (anonymous):

got it thanks.. can i throw another ques at u??

OpenStudy (dumbcow):

sure go ahead

OpenStudy (anonymous):

if x and y are the tens digit and the units digit, respectively, of the product 725,278x67,066, what is the value of x+y?

OpenStudy (anonymous):

studyn for gre..thanks

OpenStudy (dumbcow):

wait, couldn't you just multiply them together and look at the last 2 digits

OpenStudy (anonymous):

ok..i thought they would explicity state tens(which they did) and ONES digit....so units digit is equiv to ones digit?? or am i thinkn too much into it??

OpenStudy (dumbcow):

oh yeah units is same as one's place im guessing you can use a calculator

OpenStudy (anonymous):

yea they provide one..can i throw one more ques at you??

OpenStudy (dumbcow):

sure, actually the product ends up being too big, you get a 4.86X10^10 answer

OpenStudy (dumbcow):

i know for the one's digit, it will be the product of the 2 one's digits of the 2 numbers 6*8 = 48 so 8

OpenStudy (anonymous):

A developer has land that has x feet of lake frontage. The land is to be subdivided into lots, each of which is to have either 80 feet or 100 feet of lake frontage. If 1/9 of the lots are to have 80 feet of frontage each and the remaining 40 lots are to have 100 feet of frontage each, waht is the value of x?

OpenStudy (dumbcow):

hold on a sec, i'll brb

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

alright before you can find x, you need to know how many lots there are Let N be num of lots (1/9)N = N -40 N - (1/9)N = 40 8/9N = 40 N = 45 (1/9)*45 = 5 So 5 lots had 80ft, 40 had 100ft x = (5*80)+ (40*100)

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

were you able to follow that?

OpenStudy (anonymous):

a lil...what i did i got up to the 8/9x=40 but how did you get 45

OpenStudy (anonymous):

oh i got it

OpenStudy (dumbcow):

oh multiply by reciprocal

OpenStudy (anonymous):

u need hundred metals..lol thanks

OpenStudy (dumbcow):

haha thanks i didn't have to take the gre, is it a long test

OpenStudy (anonymous):

how did you get (1/9)N = N -40

OpenStudy (anonymous):

i kno im going back to soph junior yr math it sucks

OpenStudy (dumbcow):

um i got it from the statement "the remaining 40 lots" this implied that total lots(N) = (1/9)N + 40

OpenStudy (anonymous):

ok i got that

OpenStudy (dumbcow):

yeah i dunno why i just didn't write it that way lol

OpenStudy (anonymous):

no i meant n-(1/9)n=40

OpenStudy (anonymous):

o i got it

OpenStudy (dumbcow):

its all the same equation, just different ways of writing it ..thats algebra, rewriting equations

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

much thanks my test is tomorrow at 130...wish me luck..

OpenStudy (dumbcow):

good luck

OpenStudy (anonymous):

Ray is 2 inch taller than Lin, and Ray is 3 inch taller than Sam. What is the average height of ray, Lin and Sam? The median height of Ray, Lin and Sam?

OpenStudy (anonymous):

@ dumbcow...i think your answer to... if x and y are the tens digit and the units digit, respectively, of the product 725,278x67,066, what is the value of x+y? ...is wrong... ...it should be more like the answers on the yahoo link... It's kind of a trick question. You don't have to multiply a 6 digit number by a 5 digit number. If you do, you get 48641494348, and then take 4 + 8 to get 12 https://answers.yahoo.com/question/index?qid=20130603214107AAkY8t2